APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday May 21, 2006 (MJD 53877) ---=== OBSERVING TEAM ===--- Viktor Malanushenko John Barentine Shannon Watters Mike Harvanek (swing) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- A somewhat lackluster night with clouds plaguing observations throughout. We managed to finish three plates (2512, 2510, and 2204) and collect partial signal on one other (2139). ---=== OBSERVING LOG ===--- Afternoon: ---------- In anticipation of slightly warmer temperatures tonight, the spectro collimation was adjusted. The spectrograph collimation is currently set for approximately 15 C. Details are in the Problems section. Nothing else to report. Night: ------ Open at 02:05Z. Dust was high (>2000) during the second half of the night; total dust exposure was <= 8000 dust-hours. 02:40Z Cart 6 Plate 2512 (tile 1734, chunk143) GSOGTF. Seeing was good throughout, but we had light to moderate cirrus during the plate and efficiency only improved during the third exposure. Done in three exposures of 1 x 900 s and 2 x 1000 s. 04:22Z Cart 1 Plate 2510 (tile 1743, chunk143. GSOGTF. Clouds remained at bay and seeing was good. Finished in three exposures, 900, 100 and 1300 s. SoS requested a postCalib after the third exposure finished and we had sent the scope to instChange. 05:54Z Cart 3 Plate 2139 (tile 1545, chunk109) GSOGTF. This time the clouds moved in during instChange, and by the time we got set up on the field of this plate the efficiency had dropped to almost zero. We continued on because the efficiency rebounded quickly and was variable; I put us on clouds as a precaution. We aborted this exposure when the efficiency dropped solidly back to zero, and waited for the clouds to thin. After 07:00Z that was happening, so we resumed and got two exposures of 1500 s each. That was not quite enough to finish this plate before it expired at 08:35Z. 08:40Z Cart 8 Plate 2204 (tile 1588, chunk116) GSOGTF with an offset of D=-10 to bring the guide stars fully into the fibers before starting the guider. Clouds thinned during this plate to the extent that it was nearly photometric toward the end. Given the timing of our observations and when this plate expired (10:25Z, very close to the onset of twilight), we decided to squeeze in an extra exposure to get additional signal here rather than close up early. Closed for the night at 10:40Z. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6156 23:44Z 00:03Z 100 O -55.33 -50.51 46 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 39348, plate -9999 ------- 23:43 39348 0.0 bias ----- sequence 39349, plate 2204 ------- 00:05 39349 10.0 flat 00:07 39350 2.0 arc 00:27 39351 10.0 flat 00:30 39352 2.0 arc ----- sequence 39353, plate -9999 ------- 01:09 39353 0.0 bias ----- sequence 39354, plate 2512 ------- 18.0 16.8 18.9 15.2 DONE 03:13 39354 10.0 flat 03:15 39355 2.0 arc 03:38 39356 900.1 target 03:57 39357 1000.1 target 04:17 39358 1000.1 target ----- sequence 39359, plate 2510 ------- 19.8 14.9 23.7 18.0 DONE 04:28 39359 10.0 flat 04:30 39360 2.0 arc 04:49 39361 900.1 target 05:08 39362 1000.1 target 05:33 39363 1300.1 target ----- sequence 39364, plate 2139 ------- 12.3 7.0 12.0 8.9 DONE 05:51 39364 10.0 flat 05:54 39365 2.0 arc 07:51 39366 1500.1 target 08:30 39367 1500.1 target 08:33 39368 10.0 flat 08:35 39369 2.0 arc ----- sequence 39370, plate 2204 ------- 23.7 17.0 23.9 18.7 DONE 08:51 39370 10.0 flat 08:53 39371 2.0 arc 09:17 39372 1200.1 target 09:43 39373 1400.1 target 10:11 39374 1400.1 target 10:27 39375 800.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:57Z 6 0 1.88 -0.0046 74.38 0.0053 181.58 0.0405 1.000210 04:47Z 1 2510 280.27 -0.0021 63.61 0.0045 114.27 -0.0103 1.000260 07:22Z 3 2139 -75.76 -0.0098 67.37 0.0054 115.98 0.0689 1.000250 08:55Z 8 2204 -46.31 -0.0041 63.84 0.0028 141.20 -0.0146 1.000150 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 02:57Z 6 0 1.00021 -2074 -742 -150 2 74.4 17.7 2 215 -- 1.0 04:47Z 1 2510 1.00026 -2568 -1157 -175 280 63.6 17.5 6 175 -- 1.2 07:22Z 3 2139 1.00025 -2469 -987 -80 -76 67.3 13.5 8 81 -- 1.8 08:56Z 8 2204 1.00015 -1481 -233 -150 -46 63.8 15.5 3 353 -- 1.1 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:56Z 73 31 16 210 (SSW) 1221 - PtlyCldy 22:29Z 74 31 18 232 (SW) 1127 - PtlyCldy 23:01Z 75 31 14 267 (W) 1118 - PtlyCldy 23:32Z 75 30 11 203 (SSW) 1121 - PtlyCldy 00:02Z 72 28 14 231 (SW) 1162 - PtlyCldy 00:34Z 71 27 17 245 (WSW) 1176 - PtlyCldy 01:04Z 71 26 11 257 (WSW) 1185 - PtlyCldy 01:35Z 66 22 9 234 (SW) 1305 - partly cloudy 02:06Z 64 21 12 236 (SW) 1407 - PtlyCldy 02:34Z 62 21 4 244 (WSW) 1422 - PtlyCldy 03:05Z 64 21 5 197 (SSW) 1492 - PtlyCldy 03:36Z 64 21 1 178 (S) 1452 - PtlyCldy 04:08Z 64 21 5 221 (SW) 1518 - MstlyCldy 04:38Z 63 21 5 188 (S) 1539 - MstlyCldy 05:09Z 63 21 5 263 (W) 1546 - MstlyCldy 05:40Z 63 21 2 226 (SW) 1446 - MstlyCldy 06:10Z 58 30 15 79 (E) 2056 - MstlyCldy 06:42Z 57 32 7 84 (E) 1905 - MstlyCldy 07:13Z 56 33 9 69 (ENE) 2213 - MstlyCldy 07:45Z 58 30 4 89 (E) 2293 - PtlyCldy 08:15Z 58 27 6 10 (N) 2288 - PtlyCldy 08:45Z 58 27 2 77 (ENE) 2887 - PtlyCldy 09:15Z 59 28 3 312 (NW) 2883 - MstlyCldy 09:47Z 60 28 5 354 (NNW) 2710 - PtlyCldy 10:20Z 59 28 6 352 (NNW) 2436 - PtlyCldy 10:52Z 58 28 5 17 (NNE) 2303 - MstlyClr ---=== TELESCOPE STATUS ===--- 01:30Z Enclosure doors opened, fans and blowers on, louvers opened. 02:05Z Enclosure off 10:35Z Enclosure on, doors and clouvers closed, fans off Telescope parked at 121,30 with Cart 2 (Plate 2514) mounted Counterweights at 250 Spectro autofill reconnected No interlocks bypassed Dewar weights at 10:55Z Imager: 81 lbs 17 psi Spectro: 91 lbs 25 psi Note: Both dewars should last until the day crew arrives. ---=== SOFTWARE USED ===--- IOP/SOP: v4_23_0 Watcher: v2_34_0 MCP: v6_3_2 TPM: tpm_v3_3_0 AstroDa: v15_12 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_16 SoS: v5_1_0 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8530 -5.8440 Axial B -5.6850 -5.6770 Axial C 0.9410 0.9420 Trans D -10.0000 -9.9940 Lateral E 10.3505 10.3759 Lateral F 11.9380 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4711. -2457. 2997. -7825. 3296. 3024. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 22.4 Alt.: 29.997866 MIGS TONIGHT NOMINAL Axial A 1.5400 1.5420 Axial B 1.0510 1.0470 Axial C 1.1650 1.1650 Trans D -0.1240 -0.1240 GALILS Commanded: 1593419. 1603624. 1554288. -6800. -5900. Actual: 1593447. 1603642. 1554316. -6820. -5882. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1256.98 15.87 0.00 -49.55 136.74 ---=== PROBLEMS IN DETAIL ===--- Spectro Collimation Details: ---------------------------- Before collimation: sp1 Coll_motor_A 15570 Coll_motor_B 15589 Coll_motor_C 15554 sp2 Coll_motor_A 2787 Coll_motor_B -620 Coll_motor_C 173 sp1> mechPiston 2000 sp2> mechPiston 2000 After collimation: sp1 Coll_motor_A 17569 Coll_motor_B 17588 Coll_motor_C 17554 sp2 Coll_motor_A 4784 Coll_motor_B 1382 Coll_motor_C 2173 The spectrographs are currently (05-21-06) collimated for a temp of approximately +15 degree C. Bias Percentiles ---------------- Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00039353.fit -5.6 -4.3 -3.3 2.1 7.2 8.7 10.9 sdR-b1-00039353.fit -5.6 -4.3 -3.3 2.1 7.2 8.7 10.9 sdR-b2-00039353.fit -8.0 -6.6 -4.3 1.0 7.3 8.5 10.5 sdR-b2-00039353.fit -8.0 -6.6 -4.3 1.0 7.3 8.5 10.5 sdR-r1-00039353.fit -7.1 -5.4 -4.1 1.1 6.9 8.8 9.9 sdR-r1-00039353.fit -7.1 -5.4 -4.1 1.1 6.9 8.8 9.9 sdR-r2-00039353.fit -8.1 -6.3 -4.9 1.1 6.6 8.4 10.5 sdR-r2-00039353.fit -8.1 -6.3 -4.9 1.1 6.6 8.4 10.5